OpenStudy (anonymous):

Question 9 (Multiple Choice Worth 5 points) [9.02] What are the x-intercept(s) of the graph of y + 6 = x2 − x? (−2, 0) and (3, 0) (−2, 0) and (−3, 0) (2, 0) and (−3, 0) (2, 0) and (3, 0)

OpenStudy (anonymous):

will this be A ???

OpenStudy (anonymous):

x-intercept is when y=0 0 + 6 = x2 − x lets rearrange the equation 0 = x2 − x -6 two numbers that you multiply give you -6 and when you add them is -1 that will be -3 and 2 so you will have 0=(x-3)(x+2) where x=3 and x=-2

OpenStudy (anonymous):

yes you are correct again!!

OpenStudy (anonymous):

so its A

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

ok